OpenStudy (anonymous):

Rewrite the expression sin 6t cos 5t using a product identity.

OpenStudy (anonymous):

\[ \sin x \cos y=\frac{\sin(x+y)+\sin(x-y)}2 \]

OpenStudy (anonymous):

\[ \sin (\color{red}{5t}) \cos (\color{blue}{6t})=\frac{\sin(\color{red}{5t}+\color{blue}{6t})+\sin(\color{red}{5t}-\color{blue}{6t})}2 =\frac{\sin(11t)+\sin(-t)}{2} \]